$SPAD/src/input LIB.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input generated from LibraryXmpPage
)clear all
stuff := library "/tmp/Neat.stuff"
stuff.int    := 32**2
stuff."poly" := x**2 + 1
stuff.str    := "Hello"
keys stuff
stuff.poly
stuff("poly")
)system rm -rf /tmp/Neat.stuff

———-

References^{[1] nothing}

4